### v1

# Enhancing Human-Robot Collaborative Assembly in Manufacturing Systems Using Large Language Models

Abstract
The development of human-robot collaboration has the ability to improve manufacturing system performance by leveraging the unique strengths of both humans and robots. On the shop floor, human operators contribute with their adaptability and flexibility in dynamic situations, while robots provide precision and the ability to perform repetitive tasks. However, the communication gap between human operators and robots limits the collaboration and coordination of human-robot teams in manufacturing systems. Our research presents a human-robot collaborative assembly framework that utilizes a large language model for enhancing communication in manufacturing environments. The framework facilitates human-robot communication by integrating voice commands through natural language for task management. A case study for an assembly task demonstrates the framework’s ability to process natural language inputs and address real-time assembly challenges, emphasizing adaptability to language variation and efficiency in error resolution. The results suggest that large language models have the potential to improve human-robot interaction for collaborative manufacturing assembly applications.
I Introduction
Advances in robotics technology have significantly enhanced manufacturing efficiency, resulting in cost reductions and increased productivity [1, 2]. While robots can execute rapid, accurate, and repetitive tasks that demand heavy-duty efforts in manufacturing settings, they lack the capability for adaptation and versatility of human operators [2]. Therefore, the importance of Human-Robot Collaboration (HRC) is growing as humans and robots complement each other skills and capabilities. HRC refers to the interaction and cooperation between human operators and robotic systems within shared workspace [3]. Prior works have used HRC frameworks to improve ergonomics in manufacturing settings and safe human-robot interaction without compromising productivity [4, 5]. Transitioning from tasks involving large components, such as loading, placing, and unloading, to the complex assembly of smaller components such as printed circuit boards, the collaboration between human operators and robots significantly enhances both the efficiency and safety of production lines [6, 7].
Further advancing the field of HRC in manufacturing systems by improving the interaction of human operators and robots presents significant challenges. Specifically, interaction with robots induces psychological stress and tension among operators due to language barrier [8]. In contemporary manufacturing systems, human operators require extensive pre-service training and complex code development to ensure accurate and safe production with the robots [9]. These difficulties highlight the need to develop human-robot communication systems that enable interaction between humans and robots without extensive robotics training (C1). Another challenge involves the need for greater flexibility and adaptability in human-robot interaction. As the number of interactions between humans and robots on the shop floor increases, there is a higher chance of encountering unexpected changes and errors. Thus, HRC needs to be more adaptable to real-time changes and errors during the manufacturing assembly process (C2). Additionally, the demand for adaptability and flexibility extends to the need for human-robot collaborative assembly applications to integrate advanced technologies in a human-centric design to enhance communication and usability (C3).
Large Language Models (LLMs) have recently been introduced in the AI community to enhance natural language understanding and generation capabilities. These can be extended to improve human-robot interaction in the manufacturing facility. Models such as OpenAI’s GPT-3 [10] and GPT-4 [11], have shown proficiency in processing, understanding, and communicating in natural language. The integration of LLMs facilitates natural language communication between humans and robots. Using voice interaction for this communication enhances collaboration and operator safety in dynamic workspaces. The main contributions of this work are: (1) the use of LLMs for interpreting natural language to allow operators to coordinate with robotic arms, (2) a framework for system integration of voice commands, robotic arms, and vision systems to facilitate HRC in an assembly task, enhancing operational flexibility, and (3) a dynamic adaptation approach to task variations, errors, and obstacles through human-robot communication.
The rest of the manuscript is organized as follows. Section II reviews related work and identifies gaps in HRC for manufacturing assembly. Section III describes the proposed framework that integrates LLMs with voice commands, robotic arm, and sensors. Section IV showcases a case study and evaluates the framework’s performance. Finally, Section V provides conclusions and future work.
II Background
II-A Human-Robot Collaboration in Manufacturing
Recently, a wide range of methods has been developed in the field of HRC to improve the safety and efficiency of interactions between humans and robots. For instance, Mendes et al. [12] proposed an easy-to-perform HRC operating space that is more accessible to human operators. Mendes et al. integrate task-based programming and various interaction modes, including gestures and physical contact. Another work by Meziane et al. [13] developed a system for collision-free trajectories for robots in shared workspaces, utilizing a neural network for obstacle avoidance. Furthermore, the research conducted by Zanchettin et al. [14] shows that an algorithmic approach to HRCs can predict human activity patterns to control the robot more safely and effectively, resulting in a significant reduction in task time. Thus, several studies in the field of HRC have enhanced worker productivity and safety by combining the human experience with the automation capabilities of robots. Moreover, some approaches have utilized environmental data to enhance safety and efficiency. For example, Fernandez et al. [15] developed a dual-arm robotic system with multisensor capabilities for safe and efficient collaboration, integrating gesture recognition. Additionally, Wei et al. [16] gather information on human operators and the environment using RGB-D videos and predict human intentions based on a deep-learning method.
Research efforts have introduced innovative methods, such as gesture and physical contact for interaction, and neural networks for dynamic obstacle avoidance. However, there has been limited research focusing on human-robot collaborative assembly in manufacturing systems that effectively integrates interaction between humans and robots with natural language capabilities. We aim to integrate an LLM-based approach to facilitate communication into the human-robot collaborative assembly framework leveraging the strengths of both human flexibility and robot precision within the dynamic manufacturing environment.
II-B Large Language Models for Robots
Recent research has explored the integration of LLMs for robotic applications to improve human-robot interaction and robot functionalities [17, 18, 19, 20, 21, 22]. This includes the conversion of high-level human instructions into tasks that robots execute. The research by Huang et al. [17] explored leveraging LLMs, such as GPT-3 and CODEX, to convert high-level human instructions into executable actions for robots. Similarly, Singh et al. [18] developed a method called ProgPrompt. This method generates Python code for task planning that accounts for environmental conditions and pre-programmed actions. Furthermore, Lin et al. [19] developed a framework for taking natural language instructions and then developing task and motion plans. Other works in robotics have utilized LLMs to maximize the robot’s capability to comprehend and interact with its environment to resolve errors. Mees et al [20], proposed an approach to combine natural language input with visual information based on a multi-camera system to enhance task completion ability. Zhang et al. [21] utilizes LLM to address multi-agent cooperation challenges, resolving issues related to decentralized control, raw sensory data processing, and efficient task execution in embodied environments. A framework called KnowNo [22] enhances robot autonomy and efficiency in ambiguous tasks using conformal prediction to guide robots in seeking human help when faced with uncertainty.
These research studies demonstrate how integrating LLMs into robotic applications improves human-robot interaction and robot capabilities. The methodologies involve converting human instructions into executable robot actions and accounting for environmental conditions. However, existing studies have not addressed common challenges found in the manufacturing environment, such as adapting natural language communication for different operators (e.g., adapting commands based on skills, shift time, or other metrics), matching natural language to structured task sequences found in a manufacturing environment, and managing the timing of commands (e.g., identifying when to request operator assistance in the process). We aim to bridge the gap by improving the integration of natural language for human-robot interaction for an assembly process through an LLM-based framework.
III framework
Figure 1: Human-Robot Collaborative Assembly Framework Using LLM
The proposed framework, presented in Figure 1, is developed for a human-robot collaborative assembly within the manufacturing environment. The framework is designed to facilitate interaction between a human operator and a robot for an assembly process.
In this framework, a human can initialize and provide a set of tasks $T=\{t_{1},t_{2},\ldots,t_{n}\}$ as the objectives to be achieved by the robot. The robot’s capabilities $C=\{c_{1},c_{2},\ldots,c_{l}\}$ represent the specific actions that are programmed to perform within an assembly process. If the task matches the robot’s capabilities, the robot should proceed to execute the specified task $t_{i}$. That task, $t_{i}$, is broken down into subtasks (${t_{i1},t_{i2},\ldots,t_{ik}}$) that outline the sequential steps for completion. The framework has to track the last completed subtask, denoted as $t_{i\text{c}}$, where $1\leq c\leq k$ indicates the completed process within ${t_{i1},t_{i2},\ldots,t_{ik}}$. When an error is resolved by a human operator, the robot resumes from $t_{i\text{c}+1}$, ensuring continuous task progression. To enable this human-robot collaborative assembly process, this framework is structured into two primary layers: the physical layer and the virtual layer, bridging the gap between human inputs and robot actions.
III-A Physical Layer
The physical layer facilitates the interaction between the human operator and the robot within a shared physical space based on the information and data gathered from the virtual layer. The physical consists of three primary components: the human commands, the robot actions, and the sensor data. A human command is where an operator provides voice instructions to control robot actions in executing various tasks. The robot action involves the robot performing actions based on a predefined set of $T$. The sensor data is utilized to monitor environmental conditions. This data ensures the action adapts to changes in the workspace (e.g. position, orientation, or availability of parts).
When a predefined event or error $e_{i}\in E$ is detected, the robot utilizes the communication protocol to alert the human operator. This is achieved through the LLM module, which converts $e_{i}$ into a natural language message $M_{\text{e}_{i}}(t_{i})$, communicated via text-to-speech technology. Different instances of $e_{i}$ can lead to varied message $M_{\text{e}_{i}}(t_{i})$ depending on the task $t_{i}$ error. After receiving and understanding $M_{\text{e}_{i}}(t_{i})$, the human operator resolves the corresponding $e_{i}$ and commands the robot to resume $t_{i}$.
III-B Virtual Layer
The virtual layer serves to facilitate communication in natural language. This layer stores the system’s functionality to enable interaction between human commands and robot actions. The virtual layer of the system consists of two main agents: a human agent and a robot agent.
III-B1 Human Agent
We have integrated a human agent into our system to enhance the human-robot interaction based on the approach by Zhang et al. [23]. The goal of the human agent in our proposed framework is to process the human commands in voice instructions into a text that the robot agent can understand. This agent ensures that the human command intentions are transferred to the robot agent without the need for manual programming. The human agent contains an Automatic Speech Recognition (ASR) module responsible for processing the human command voice data, which converts this data into text that can be processed by the robot agent. The communication module in the human agent serves as a bridge, transmitting commands and information to the robot.
III-B2 Robot Agent
The robot agent interprets and executes the set of manufacturing tasks $T$ based on the voice commands received from the human operator. This process is facilitated by various functional modules:
  • Initialization Module: The initialization of the robot agent, provides basic operational guidelines and task execution protocols. This process involves programming the robot with instructions to perform tasks $T$ within its capabilities $C$ and to communicate errors or request assistance from human operators when necessary. The following initial prompt is given to the robot agent:
    "You are a robot agent in a human-robot collaborative assembly system designed to assist in tasks and respond to commands. Upon receiving a request within your capability range, execute the service. In the event of encountering errors, request assistance from a human operator for error correction, providing clear and understandable explanations."
    This initialization prompt defines the robot agent’s role in responding to commands. It also establishes a protocol for seeking assistance and communicating errors to human operators for task completion. Following the initial setup of the robot agent, specific capabilities $C$ are defined, such as specific assembly tasks.
  • LLM Module: A LLM is used to interpret human commands into a corresponding task $t_{i}\in T$. Human commands and functional capabilities $C$ including the detailed functional information of $C$ are put into the LLM. The LLM then analyzes the language and context of the human command. This process allows the system to match the command with the predefined functional capabilities $C$, ensuring the intended task is accurately executed. The LLM module also communicates any detected issues from the task control module into understandable natural language for the human operators, facilitating HRC for error resolution.
  • Sensor Module: The sensor module processes data from the physical layer such as position, orientation, and part availability. This information is utilized for adjusting robot actions required for executing manufacturing tasks $T$. For example, the module provides component positions and orientations for the task control module. This enables accurate robotic adjustments for assembly or initiates error notifications for missing components, ensuring effective task execution and error management.
  • Task Control Module: The task control module directs robot actions to fulfill task $t_{i}$ within the set $T$, based on the functional capabilities $C$. The task control module adjusts robot actions to environmental conditions obtained from the sensor module. Additionally, this module plays a role in managing errors by verifying sensor data and communicating detected issues through the LLM module to facilitate resolution by the human operator.
III-C Human-Robot Collaborative Assembly Workflow
The overall workflow is depicted in a sequence diagram,  Figure 2, which visually represents the process flow within a human-robot collaborative assembly framework. The sequence diagram shows how voice commands from the human operator are processed and interpreted by the LLM Module and guides the robot actions.
The process begins with a human operator delivering a voice command, which is converted into a discrete set of tasks $T$ by the LLM module that the robot agent is programmed to understand and execute. The robot then requests sensor data to execute $t_{i}$. If the data provided by the sensor is valid, the robot proceeds to execute the assigned $t_{i}$. The sensor module determines data validity by comparing detected parameters to predefined criteria. A successful task execution results in a completion message $M_{\text{c}}(t_{i})$ to the human operator via the LLM Module.
When sensor data is invalid, or $t_{i}$ has any errors, the robot agent initiates the communication protocol to generate an error message $M_{\text{e}_{i}}(t_{i})$ via the LLM module, aiming to inform the human operator of the specific error and its occurrence within subtask $t_{i\text{c}+1}$ for efficient resolution. Following the error identification and correction by the human operator, a new command by the human operator is issued to the robot. The robot then resumes task execution at $t_{i}$, starting from the interrupted subtask $t_{i\text{c}+1}$, based on the new sensor data. This procedure is repeated until $t_{i}$ is successfully completed.
IV Case Study
Figure 2: Sequence Diagram for Human-Robot Collaborative Assembly in Manufacturing Systems
The proposed framework was tested in a manufacturing assembly manufacturing cell. The goal of the manufacturing cell is to assemble a cable shark product. The cable shark assembly contains (1) a housing, (2) a wedge, (3) a spring, (4) and an end cap. The exploded cable shark assembly and the final assembled product are shown in Figure 3. This assembly process task $t_{1}$ include four sequential subtasks: housing assembly ${t_{11}}$, wedge assembly ${t_{12}}$, spring assembly ${t_{13}}$, and end cap assembly ${t_{14}}$.
The physical setup is implemented using a Six Degrees of Freedom (6DOF) UFactory xArm with a UFactory 2-finger gripper as the end effector [24]. The base of the robotic arm remains fixed at the center, accompanied by a mat on one side, while the parts are assembled on the adjacent side. Two fixed cameras as visual sensors were placed in the surrounding area, one focused on the mat area for pre-assembly component placement, and the other camera on the assembling process. The conversion between camera coordinates and robot arm coordinates is accomplished through the coordinate transformation, as described in Section IV-C.
IV-A LLM and ASR Module
This section outlines how LLM and the ASR module, were implemented within the system. The communication aspect within the system is enabled by OpenAI’s speech-to-text and text-to-speech models. The OpenAI’s transcription model ‘whisper1’ [25] is implemented to transform speech-to-text, ensuring that human voice instructions are accurately captured. To enable the communication protocol to provide a verbal response from the robot, OpenAI’s text-to-speech model ‘tts-1’ [26] is utilized to generate audio responses. This approach enhances and allows for a clear exchange of information between the human and the robot.
The LLM module, enabled by OpenAI’s GPT-4.0 [11], is the decision-making component of the proposed framework. The LLM module converts instructions from the human operator into tasks $T$ executable by the robot, utilizing its capabilities $C$ through OpenAI’s function calling feature [27]. This module ensures that the robot actions match the human operator commands, allowing for adaptive and responsive task management within the manufacturing environment.
Figure 3: Cable Shark Assembly
IV-B Sensor Module: Vision System
Figure 4: Feature Extraction Method with the Vision System
We incorporated a vision system as our sensor module. During the collaborative assembly process, the vision system provides feedback on environmental data to the task control module. This enables precise object detection and identification of object orientation, facilitating task execution and error management. To enable object detection, Yolov5, a computer vision model, is utilized [28]. Custom Yolov5 models are trained using a dataset assembled for object detection. This dataset comprises images of individual parts (i.e. housing, wedge, spring, end cap), with manual bounding box annotations around each part in the images, and distinct classes for each part are created. The vision system utilizes this dataset with bounded boxes as a reference to identify the parts, as shown in Figure 4. If a part is detected, the coordinates of the top-left and bottom-right corners of the bounding box are retrievable. Once these coordinates are obtained, the x and y coordinates of these points are utilized to determine the object’s midpoint. This midpoint serves as the target point for the robotic arm to pick up the part. Following the successful object identification process, the valid sensor data is transferred to the task control module for guiding robot action. For parts that are not accurately detected, the sensor module transfers invalid data to the task control module to notify the human operator via LLM.
IV-C Task Control Module: Assembly Task
Figure 5: Cable Shark Assembly Process
TABLE I: Language Variations for Task Instructions
Scenario 1 Instruction Type Instruction
Scenario 1: Component Overlap Specific ["Overlap resolved. Proceed with the task."], ["Problem is…"]
Moderately Specific ["I’ve placed the components correctly."], ["I’ve sorted out the…"]
Least Specific ["Fixed."], ["Done."], ["Completed."], ["Handled."], ["Adjusted."]
Scenario 2: Incorrectly Assembled Part Specific ["Correction is made. Resume the task."], ["The wedge is…"]
Moderately Specific ["I’ve fixed the issue with the wedge."], ["I’ve placed the wedge…"]
Least Specific ["Fixed."], ["Done."], ["Addressed."], ["All set."], ["Under control."]
Scenario 3: Missing Component Specific ["I’ve placed the spring component. Please proceed."].["Spring…"]
Moderately Specific ["I’ve fixed the issue with the spring."], ["The spring component…"]
Least Specific ["Fixed."], ["Done."], ["Managed."], ["Handled"], ["Settled."]
TABLE II: Success Rates for Language Variations
Scenario Instruction Category Success Rate
Scenario 1: Component Overlap Specific 100%
Moderately Specific 73%
Least Specific 27%
Scenario 2: Incorrectly Assembled Part Specific 93%
Moderately Specific 87%
Least Specific 53%
Scenario 3: Missing Component Specific 100%
Moderately Specific 67%
Least Specific 27%
The objective of the Task Control Module is to execute tasks as interpreted from human commands by LLM and to manage errors within a human-robot collaborative assembly system. Error handling is also a critical function of this module. The task control module verifies sensor data from the vision system. For instance, this module checks if a component is properly positioned for a pickup. If the sensor data is valid, the task control module continues with its operation. When the sensor data is invalid (e.g. missing component), the module transfers detailed error information to the LLM module to convert the information into natural language. This process activates the communication protocol, allowing the human operator to recognize and address the error.
To enable the assembly process, the task control module integrates various functionalities, including coordinate transformation and position and rotation control, based on the sensor data obtained from the vision system. These functions enable the execution of tasks interpreted by the LLM. The coordinate transformation is conducted to align the robotic arm ($A$) and camera ($C$) coordinate systems within the assembly process’s base frame ($B$) to ensure an accurate assembly task. The transformation between the robot’s base (${}_{B}^{A}p$) and camera’s (${}_{B}^{C}p$) positions is defined by
${}_{B}^{A}p=_{C}^{A}T\cdot_{B}^{C}p,$ (1)
where ${}_{C}^{A}T$ is the transformation matrix incorporating translation and rotation. The matrix includes a 3x3 rotation matrix ${}_{C}^{A}R$ and a translation vector $t$, translating coordinates from the camera to the robotic arm system. The effective calculation of positions $(x,y)$ relative to the base frame relies on this transformation for the robotic arm’s precise movements and interactions within the assembly setup. The position and rotation control function is then implemented to fine-tune both the positional alignment and the orientation of components for assembly tasks. To translate camera-detected object coordinates ($x_{p},y_{p}$) from pixel values to the world coordinate system, We used predefined offset values ($x_{o},y_{o}$), calculated as
$V=x_{o}\times x_{p},\quad H=y_{o}\times y_{p},$ (2)
where $V$ and $H$ adjust the arm’s trajectory based on the visual sensor input. Concurrently, the rotation control mechanism assigns specific trajectory angles to the robotic gripper, ensuring precise orientation alignment of parts for insertion. The cable shark assembly process is shown in Figure 5.
IV-D Case Study Results
The proposed framework was integrated into the assembly system to study the effect of integrating LLMs with a knowledgeable operator [^1] See supplementary video file for a demonstration of the human-robot collaborative assembly.. The operator knew the assembly process and was able to converse. Three scenarios were used to evaluate the proposed framework:
Scenario 1: The system detects an overlap of the housing components and requests human intervention.
Scenario 2: If a wedge component is incorrectly assembled, the robot halts and requests human correction.
Scenario 3: A missing spring component is detected and the robot requests the human operator to place the spring component.
Based on these scenarios, the evaluation focused on assessing the system’s proficiency in interpreting and performing tasks derived from human commands, particularly considering the linguistic and human diversity. The variations in phrasing and terminology across different commands were cataloged, as presented in Table I, to validate the system’s adaptability to diverse linguistic inputs. The instruction type was classified into three distinct categories: specific, moderately specific, and least specific. Five unique variations of each instruction type were provided across three distinct tasks. To ensure the reliability of the evaluation, each language variation scenario presented in Table I is conducted with three repetitions.
Figure 6: Case Study Communication Results for Each Scenario
The results highlight the effectiveness of integrating LLM into the human-robot collaborative assembly framework. Figure 6 showcases the human-robot communication utilizing the vision system, outlined in Section IV-B, to ensure task completion in each scenario. In scenario 1, the robot detects overlapping components which signal to human operators for help. Upon resolution, the human operator prompts task continuation. Similarly in scenario 2, the robot identifies a misassembled wedge, notifying the human operator, and then the human commands the robot to continue after manually assembling the wedge. For Scenario 3, the robot flags a missing spring component, requesting human intervention, and the human instructs the robot to resume the assembly once the issue is fixed.
Based on the evaluation setup, we assessed the system’s capability to understand and execute commands with varied language expressions. For tasks with specific instructions, success rates were high, averaging approximately 98%. The rates decreased to 76% with moderately specific instructions and dropped further to 36% with least specific instructions. The success rates for each category of instruction type are detailed in Table II. This data suggests a positive correlation between instruction type and task execution success.
IV-E Case Study Discussion and Limitations
The case study evaluated the system’s ability to facilitate collaboration between humans and robots and showcases how the integration of LLMs can lead to efficient and flexible manufacturing processes. The results demonstrate that as the instructions become less specific, the robot’s performance decreases significantly, indicating that the robot agent needs well-defined human commands to operate effectively. These results highlight the limitations of the proposed framework and suggest an area for improvement. One such improvement involves introducing feedback mechanisms that allow the system to ask for clarification when instructions are unclear, refining the interaction between humans and robots. Notably, the robot’s success was higher with commands that included keywords like ’fixed’ or ’addressed,’ which means that the robot understands its initialization. The initialization protocol informs the robot that the human’s role is to correct errors to resume the assembly task. Based on the case study, we have found that the initialization process is crucial as it ensures that the robot can understand the roles within the system.
There were also limitations in the case study and evaluation of the proposed framework. Specifically, we only evaluated a limited range of commands from the operator. These commands were specifically related to predefined assembly scenarios and did not include other interactions, such as human interruptions and task-irrelevant questions from the human. The case study also did not study the variability in the operator’s knowledge, as the framework assumed operators would provide relevant instructions to the assembly task. Furthermore, the operator was not allowed to change the defined tasks, e.g., variations in positioning, or task order. Future work will focus on analyzing the developed framework in a variety of manufacturing scenarios. We will look to test the proposed system with operators who do not have prior knowledge of the assembly process. We will also test the proposed framework for other material handling and manufacturing tasks.
V Conclusion and Future Work
Advancements in LLM are applied to human-robot collaborative assembly for executing robot actions and collaborating with humans based on environmental data. By incorporating LLM, robots can better understand and interact with human operators to resolve errors and improve execution based on environmental feedback. Drawing on insights from these studies, we incorporated LLM into our human-robot collaborative assembly framework, enabling dynamic responses to task variations.
This research addresses key challenges in human-robot collaborative assembly, including developing communication systems that require minimal robotics training (C1), enhancing adaptability and flexibility to manage real-time changes and errors in human-robot collaborative assembly within the manufacturing environment (C2), and integrating advanced technologies into human-robot collaborative assembly process with a human-centric design to improve usability and communication (C3). To validate the effectiveness of our proposed framework, we implemented a practical setup involving the cable shark device assembly process. This setup showcased the framework’s capability to facilitate intuitive human-robot communication through voice commands with language variations. We dynamically adapt to task variations and errors by integrating LLM, sensors, and task control mechanisms, demonstrating its ability to maintain productivity and ensure a continuous workflow.
Future work includes extending this LLM-based human-robot collaborative assembly framework by focusing on increasing its adaptability by feeding LLM with a wide range of data, such as various robotic tasks and sensor data. The aim is to increase the robot’s flexibility in performing tasks, enhance safety, and enable it to address unexpected errors and task variations. Additionally, we also plan to incorporate multiple modalities, such as haptic and gestures to improve human-robot interaction within the manufacturing environment. Our framework plans to enhance this multimodal strategy by integrating LLM to improve communication efficiency, task adaptability, and intuitive interaction between humans and robots within the manufacturing environment.
Acknowledgment
We thank Dana Smith from DMI Companies, Inc. for providing the case study and offering valuable feedback on the project.